# LLaVA-Gemma: Accelerating Multimodal Foundation Models with a Compact Language Model

### v1

@@ 3.4 Speed of Training and Inference
We compare the training and eval speed for the two models sizes. In our experiments, the training time for the Gemma-2B model on 8 Intel Gaudi 2® AI accelerators was 4 hours, while the larger Gemma-7B model required 16 hours to train under the same conditions. This indicates that the Gemma-7B model, with its increased parameter count, takes approximately four times longer to train compared to the Gemma-2B model. The relative speed of the Gemma-7B model is thus 0.25x compared to the Gemma-2B model. These results highlight the trade-off between model size and training efficiency, with larger models requiring significantly more computational resources and time.

### v2

@@ 3 Results
One item of note is that for the ScienceQA dataset, the larger models consistently perform better than smaller due to the datasets task requiring diverse general knowledge captured better by the larger models.
@@ 3.4 Speed of Training and Inference
We compare the training and eval speed for the two models sizes. In our experiments, the training time for the Gemma-2B model on 8 Intel Gaudi 2® AI accelerators was 4 hours, while the larger Gemma-7B model required 16 hours to train under the same conditions. This indicates that the Gemma-7B model, with its increased parameter count, takes approximately four times longer to train compared to the Gemma-2B model. The relative speed of the Gemma-7B model is thus 0.25x compared to the Gemma-2B model. We find a similar speed ratio during inference. These results highlight the trade-off between model size and training efficiency, with larger models requiring significantly more computational resources.